lolol:

Assume f(x) = -2x + 8 and g(x) = 3x. What is the value of (f + g)(-1)?

psirockin2:

so you want (f+g)(-1) (-2x+8+3x)(-1) So that'd be -x-8

wo1f0mon:

correct btw remember P.E.M.D.A.S.

Nnesha:

hmm you're not multiplying (f+g) with (-1)

Nnesha:

(f+g)(-1) means add both function and then substitute x for -1

Nnesha:

\[f(x)=-2x+8 ~~~~g(x)=3x\]\[f(x)+g(x) =(-2x+8)+(3x)\] combine like terms add the coefficient of the same variables like `-2x+3x` (-2+3)x=1x\[f(x)+g(x)=3x-2x+8= x+8\] now replace x with `-1` and then simplify that would be the answer